a car salesperson as arrange a visit on three perspective customers in a week based on the past experiences helpers and laws that there is 10% chance of closing a sale on each visit determine the probability distribution of the number of sales and persons will make


1
Expert's answer
2022-02-21T17:55:33-0500

This is a binomial distribution with "n=3,p=0.1."

"P(X=0)=C_3^00.1^00.9^3=0.729."

"P(X=1)=C_3^10.1^10.9^2=0.243."

"P(X=2)=C_3^20.1^20.9^1=0.027."

"P(X=3)=C_3^30.1^30.9^0=0.001."